Maranello Inc. sells ceramic mugs. During May 2017, its inventory records for one brand of ceramic mugs were as follows:

Beginning Inventory ....................

20 mugs

@

$ 3

May 1 Purchase .....................

24 mugs

@

3.25

May 3 Sale ........................

35 mugs

May 7 Purchase ....................

57 mugs

@

3.50

May 9 Sale ........................

30 mugs

  1. See information for Maranello Inc. above. Using this information, periodic FIFO cost of goods sold is
  1. $211.50
  2. $225.50
  3. $153.75
  4. None of the above
  1. See information for Maranello Inc. above. Using this information, the cost of goods sold using the perpetual average cost method is (round your answer to two decimals)
  1. $213.29
  2. $217.20
  3. $214.77
  4. None of the above
Answer #1
1
FIFO cost of goods sold
May 3 Sale:
From Beginning Inventory 60 =20*3
From May 1 Purchase 48.75 =(35-20)*3.25
May 3 Sale Cost of goods sold 108.75
May 9 Sale:
From May 1 Purchase 29.25 =(9*3.25)
May 7 Purchase 73.50 =(30-9)*3.50
May 9 Sale Cost of goods sold 102.75
FIFO cost of goods sold 211.50 =108.75+102.75
Option A $211.50 is correct
2
Average cost of May 3 Sale 3.136 =((20*3)+(24*3.25))/(20+24)
May 3 Sale Cost of goods sold 109.78 =35*3.136
Average cost of May 9 Sale 3.450 =((9*3.136)+(57*3.50))/(9+57)
May 9 Sale Cost of goods sold 103.51 =30*3.450
Perpetual average cost of goods sold 213.29 =109.78+103.51
Option A $213.29 is correct